OpenStudy (anonymous):

how to integrate xe^-x^2 with integration substitution?

OpenStudy (anonymous):

try \(u=-x^2,du=-2xdx,-\frac{1}{2}du=xdx\) and you are done in pretty much one step

OpenStudy (anonymous):

you get \[-\frac{1}{2}\int e^udu\]

OpenStudy (anonymous):

oh right! thank you!

OpenStudy (anonymous):

and since \(e^u\) is its own anti derivative, you ge t \[-\frac{1}{2}e^{-x^2}\] soon you will be able to do these in your head

OpenStudy (anonymous):

yw